In a certain county 35% of the people have more than one sibling. Suppose a simple random sample of 650 people is selected. Use the central limit theorem to calculate the probability that less than 37% of these people have more than one sibling.
Answer #1
for normal distribution z score =(p̂-p)/σp
here population proportion=     p= 0.350
sample size       =n= 650
std error of proportion=σp=√(p*(1-p)/n)= 0.0187
probability = P(X<0.37) = P(Z<1.07)= 0.8577

(try 0.8575 if this comes wrong and revert)